Calculate the molar mass of the following compounds. Fe2O3 and H2SO4 (Al:27, Fe:56, S:32, O:16, H:1)


1
Expert's answer
2021-04-01T06:01:58-0400

"Fe_2O_3=(56\times2)+(16\times3)"

"=112+48=160"

"H_2SO_4=(1\times2)+(32)+(16\times4)"

"=2+32+64=98"